OpenStudy (anonymous):

lost yet again -4/5x=3/5

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I got -12/25 but this does not work when i plug it in

OpenStudy (lgbasallote):

is that \[\huge -\frac 45 x = \frac 35\]

OpenStudy (anonymous):

yes, it is how did you type in the fraction?

OpenStudy (lgbasallote):

try clicking the equation button by the comment box...you should see the fraction command there

OpenStudy (lgbasallote):

anyway if you multiply both sides by 5 what do you get?

OpenStudy (anonymous):

ok i see it now thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

for the -4/5 * 5 =-4 and 3/5*5=3

OpenStudy (lgbasallote):

yes dont forget x

OpenStudy (anonymous):

x=-4/3

OpenStudy (anonymous):

wow I get so lost so quick

OpenStudy (lgbasallote):

it's not -4/3

OpenStudy (anonymous):

y, what did i do wrong

OpenStudy (lgbasallote):

i assume you tried to solve it mentally. that's what you did wrong. do it step by step

OpenStudy (anonymous):

ok give me a min

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

do i flip the -4/5 when I isolate the x like this -4/5x=3/5 x= 3/5 div -5/4

OpenStudy (lgbasallote):

\[\huge -\frac 45 x = \frac 35\] you start by multiplying both sides by 5

OpenStudy (anonymous):

y 5

OpenStudy (anonymous):

is it because both of the fraction have 5?

OpenStudy (lgbasallote):

you multiply both sides by 5 because x has a denominator of 5

OpenStudy (anonymous):

ok so this gets rid of the fractions

OpenStudy (anonymous):

right ?

OpenStudy (lgbasallote):

yes. so what does the equation look like after?

OpenStudy (anonymous):

-4x=3 ?

OpenStudy (lgbasallote):

right. so now what would you do?

OpenStudy (anonymous):

div both sides by -4

OpenStudy (anonymous):

-0.75

OpenStudy (anonymous):

x=0.75

OpenStudy (anonymous):

-0.75 my bad

OpenStudy (lgbasallote):

yes that is right. but you should but it in fraction for

OpenStudy (anonymous):

-4/3

OpenStudy (lgbasallote):

form*

OpenStudy (lgbasallote):

how did you get -4/3? (after i said it wasnt right)

OpenStudy (anonymous):

1 1/3 sorry

OpenStudy (anonymous):

looked at the wrong part of my paper

OpenStudy (anonymous):

1/3

OpenStudy (lgbasallote):

....what did you get again after you divided by 5?

OpenStudy (anonymous):

-4x=3

OpenStudy (lgbasallote):

so what did you after that?

OpenStudy (anonymous):

-4x/5=3/5 =>-4x=3*5/5 =>-4x=3 =>x=-3/4

OpenStudy (lgbasallote):

what did you do*

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

wow I have been typing it upside down

OpenStudy (lgbasallote):

what were you doing to get -4/3?

OpenStudy (anonymous):

forgot to div both sides by -4 and because i can not reduce -3/4 that is my answer

OpenStudy (lgbasallote):

you need to remember that you divide both sides by the coefficient of x ALWAYS... in this case your coefficient is 4..that's why you divide both sides by 4

OpenStudy (anonymous):

because what i do to one side I do to the other

OpenStudy (anonymous):

in order to get x alone I div -4x by -4 and I div 3 by -4 also this gives me -3/4

OpenStudy (anonymous):

guys this problem is simple -4/5x=3/5 take the -4/5 and divide by 3/5 so you have X= -4/5 divided by 3/5 now flip the 3/5 to 5/3 so that is to change the divide to times so now you have -4/5 times 5/3= ??

OpenStudy (anonymous):

-4/3

OpenStudy (anonymous):

I am so wrong just looked at it

OpenStudy (lgbasallote):

uhhh what?

OpenStudy (lgbasallote):

so what's your answer...-3/4 or -4/3?

OpenStudy (anonymous):

IDK!!!!!!!!

OpenStudy (anonymous):

sorry for yelling

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

you have -4x = 3 you said you divide by =4 so what happens if you do that?

OpenStudy (anonymous):

the web site wolfram alfa said it is -4/3

OpenStudy (lgbasallote):

could you post the URL of that wolframalpha?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-4%2F5x%3D3%2F5 here it is

OpenStudy (anonymous):

no need that is right -4/3 that is what I got when I worked the problem out as well.look at my post where I went step by step through this problem.

OpenStudy (anonymous):

reduced form is 1and 1/3

OpenStudy (anonymous):

nooo

OpenStudy (lgbasallote):

you told be the problem was \[\huge -\frac 45 x = \frac 35\] what you posted in wolframalpha was \[\huge -\frac{4}{5x} = \frac 35\] see the difference/

OpenStudy (lgbasallote):

you told me*

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

in the first one, the x is in the numerator. In the second one, the x is in the denominator

OpenStudy (lgbasallote):

so which one is it?

OpenStudy (anonymous):

what is this again??

OpenStudy (anonymous):

the x is off to the side of the side like -4/4(x)

OpenStudy (lgbasallote):

then like i said -4/3 is wrong

OpenStudy (lgbasallote):

this is how you type it in wolfram http://www.wolframalpha.com/input/?i=-4%2F5+x%3D3%2F5

OpenStudy (lgbasallote):

wolfram is very meticulous and choosy so you have to wrie the question properly

OpenStudy (lgbasallote):

got it now?

OpenStudy (lgbasallote):

@tootie149 are you still there?

OpenStudy (theviper):

\[\huge -\frac 45 x = \frac 35\] \(\LARGE{-4x=\frac{3\times5}{5}}\) \(\LARGE{-4x=3}\) \(\LARGE{x=\frac{3}{-4}}\)